OpenStudy (delaneyparris):

what is b in 1/2(-3)=5/2

OpenStudy (anonymous):

Distribute the 1/2 1/2b - 3/2 = 5/2 1/2b = 5/2 + 3/2 1/2b = 8/2 1/2b(2) = 8/2(2) b = 8